# Tensor Attention Training: Provably Efficient Learning of Higher-order Transformers

### v1

@@
Jiuxiang Gu jigu@adobe.com. Adobe Research.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song zsong@adobe.com. Adobe Research.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
Tensor Attention, a multi-view attention that is able to capture high-order correlations among multiple modalities, can overcome the representational limitations of classical matrix attention. However, the $\Omega(n^{3})$ time complexity of tensor attention poses a significant obstacle to its practical implementation in transformers, where $n$ is the input sequence length. In this work, we prove that the backward gradient of tensor attention training can be computed in almost linear $n^{1+o(1)}$ time, the same complexity as its forward computation under a bounded entries assumption. We provide a closed-form solution for the gradient and propose a fast computation method utilizing polynomial approximation methods and tensor algebraic tricks. Furthermore, we prove the necessity and tightness of our assumption through hardness analysis, showing that slightly weakening it renders the gradient problem unsolvable in truly subcubic time. Our theoretical results establish the feasibility of efficient higher-order transformer training and may facilitate practical applications of tensor attention architectures.
@@ 1 Introduction
The generative large language models (LLMs), such as Mistral [65], Llama [121], Llama2 [123], Llama3 [15], Gemma [122], GPT-3 [28], GPT-4 [2], Claude3 [18], Grok-1 [136] and many more have been widely involved in people’s living and work in these two years, such as bio-informatics [124], coding [60], education [70], finance [83], law [117], medicine [125], and even writing NeurIPS conference reviews [76]. The success of LLMs is based on the transformer architecture introduced by [128], which also has been introduced into other modality [41], such as vision-language models, e.g., CLIP [100], Flamingo [8], LLaMA-Adapter [149, 50], LLava [80, 77], BLIP [81, 79], MiniGPT-4 [144], Qwen [23, 24], Gemini [119], MM1 [87]. It is worth noting that all the above open-sourced large models use two-view matrix attention, i.e., each attention score/entry is related to two tokens (one query token and one key token) to capture the data correlation. However, nowadays, many studies find that multi-view is crucial for high-order correlation in various kinds of data, e.g., math [112], graph [43, 84], and multi-modality [74].
Recently, OpenAI released GPT-4o [95], and Google released Project Astra [54], two flagship multi-modality models that can reason across three views, i.e., audio, vision, and text in real-time, shocking the community again. One fundamental technical obstacle for multi-modality models is efficiently fusing multiple representations/views to capture the high-order correlation among different modalities. More specifically, let $Z$ be hidden representations and $Q,K,V$ be corresponding query, key, and value matrices. Then, the classical/matrix attention head can be written as $\mathsf{Att}(Z)=\mathsf{softmax}(QK^{\top})V$. Denote $A,B,C$ as three modalities. The three mainly existing fusion methods are (Table 2 in [139]): (1) Summation, $\mathsf{Att}(Z_{A}+Z_{B}+Z_{C})$; (2) Concat., $\mathsf{Att}([Z_{A},Z_{B},Z_{C}])$; (3) Multiple cross attentions, $\mathsf{softmax}(Q_{A}K_{B}^{\top})V_{B}$. However, [112] theoretically and empirically shows that all three methods can only capture pairwise correlation but not triple-wise correlation due to the representational limitations of matrix attention. In other words, one classical matrix attention head “cannot” capture the information relevant to the image, audio, and text simultaneously unless using multiple layers with careful architecture design.
@@
There are several brilliant latest previous works to overcome the time complexity bottleneck above, e.g., $\Omega(n^{2})$ for matrix attention and $\Omega(n^{3})$ for tensor attention. [148] accelerate matrix attention forward via kernel density estimation and get truly sub-quadratic time running time. [19] uses the polynomial approximation method to map the matrix attention into low-rank matrices during forward computation, leading to an almost linear time complexity $n^{1+o(1)}$ when entries are bounded. Similarly, under sparsity assumptions, [56] achieves nearly linear time computation for matrix attention forward by identifying the larger entries in the attention matrix. On the one hand, with fine-grained analysis, [20] proposes a new backward algorithm to compute the gradient of matrix attention in almost linear time complexity $n^{1+o(1)}$ as well, under the same bounded entry assumption. On the other hand, [21] surprisingly finds that the forward computation of tensor attention can also be achieved in almost linear time $n^{1+o(1)}$ rather than almost quadratic time $n^{2+o(1)}$, under similar assumptions as [19]. See a summary in Table 1. Thus, it is natural to ask,
Can we achieve almost linear time complexity during Tensor Attention Training?
Figure 1: The visualization of tensor attention with $\mathsf{softmax}$ activation function (Definition 3.7). We give an example of input token length $n=8$, feature dimension $d=4$. Here, both $\mathsf{X},\mathsf{Y}$ are tenors in shape $d\times d\times d$. There are three input matrices for each tensor.
The answer is positive. In this work, under the same bounded entries assumption as [21], we propose a method to fast compute the backward gradient of Tensor Attention Training in almost linear time as its forward computation. Thus, our results may make the tensor attention practical, as we can get around the $\Omega(n^{3})$ complexity barrier both in its forward and backward computation. We summarize our contribution as below:
Our contribution:
@@
  • Based on the closed-form solution, by utilizing polynomial approximation methods and tensor computation tricks, we propose a method to fast compute the backward gradient of tensor attention training in almost linear time as its forward computation (Theorem 5.2).
@@ 2 Related Work
In recent years, significant advances have been made in the development of efficient attention computation. One research direction involves employing low-rank approximations, polynomial kernel, or random features for the attention matrix [103, 78, 134, 32, 33, 151, 1, 19, 68, 110, 5], which scales the computational complexity linearly with sequence length. Another method explores patterns of sparse attention that lessen the computational load [31, 30, 147, 108, 56, 53]. Additionally, using linear attention as an alternative to softmax attention has emerged as a substantial area of study [120, 71, 114, 146, 109, 5, 118, 143, 51, 45]. These innovations have enhanced the capability of transformer-based models to handle longer sequences, thereby broadening their potential applications across various fields [37, 105, 99, 46, 92, 138, 14, 22, 62].
@@
Tensors excel over matrices in capturing higher-order relationships within data. Calculating low-rank factorizations or approximations of tensors is essential in a wide range of computer science applications, such as natural language processing [38, 85, 86, 29], computer vision [129, 131, 111, 58, 55, 9, 98, 75, 34], computer graphics [130, 135, 127], security [4, 7, 67], cryptography [47, 107, 113, 73], computational biology [39, 106], and data mining [66, 102, 69, 90]. Moreover, tensors are crucial in numerous machine learning applications [91, 10, 97, 63, 57, 17, 3, 11, 12, 26, 49, 64, 16, 13, 150, 142, 115] and other diverse fields [104, 132, 42, 35, 36, 140, 116, 82, 96, 145, 141, 101].
Roadmap.
In Section 3, we introduce the notations we use, several useful definitions and our loss function. In Section 4, we give the closed form of the gradient of our loss function, and also its computational time complexity. In Section 5, we prove that we can compute the gradient in almost linear time. In Section 6, we analyze the hardness of our algorithm. In Section 7, we give the conclusion of our paper.
@@ 3 Preliminary
In this section, we first provide the notations we use. In Section 3.1, we provide general definitions related to tensor operation. In Section 3.2, we provide key definitions that we will utilize in this paper.
@@
Basic notations.
We use $[n]$ to denote $\{1,2,\dots,n\}$. We use $e_{i}$ to denote a column vector where only $i$-th location is $1$ and zeros everywhere else. We denote an all $1$ vector using ${\bf 1}_{n}\in\mathbb{R}^{n}$ We use $\langle a,b\rangle$ to denote the inner product of $a,b\in\mathbb{R}^{d}$ i.e. $\langle a,b\rangle:=\sum_{i=1}^{d}a_{i}b_{i}$. We use $\|x\|_{p}$ to denote the $\ell_{p}$ norm of a vector $x\in\mathbb{R}^{n}$, i.e. $\|x\|_{2}:=(\sum_{i=1}^{n}x_{i}^{2})^{1/2}$, and $\|x\|_{\infty}:=\max_{i\in[n]}|x_{i}|$. We use $\operatorname{tr}[A]:=\sum_{i=1}^{n}A_{i,i}$ to denote the trace of a matrix $A\in\mathbb{R}^{n\times n}$. We use $\exp(A)$ to denote a matrix where $\exp(A)_{i,j}:=\exp(A_{i,j})$ for a matrix $A\in\mathbb{R}^{n\times d}$. We use $\|A\|_{\infty}$ to denote the $\ell_{\infty}$ norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e. $\|A\|_{\infty}:=\max_{i\in[n],j\in[d]}|A_{i,j}|$. We use $\|A\|_{F}$ to denote the Frobenius norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e. $\|A\|_{F}:=\sqrt{\sum_{i\in[n]}\sum_{j\in[d]}|A_{i,j}|^{2}}$. We use $\operatorname{poly}(n)$ to denote polynomial time complexity w.r.t $n$.
Tensor related notations.
Let $A\in\mathbb{R}^{n\times d}$. We use $a:=\operatorname{vec}(A)$ to denote a length $nd$ vector. We stack rows of $A$ into a column vector, i.e. $\operatorname{vec}(A):=[a_{1}^{\top},a_{2}^{\top},\dots,a_{n}^{\top}]^{\top}$ where $a_{i}^{\top}$ is the $i$-th row of $A$, or simply $\operatorname{vec}(A)_{j+(i-1)d}:=A_{i,j}$ for any $i\in[n],j\in[d]$. Let $\mathsf{A}\in\mathbb{R}^{d_{1}\times d_{2}\times d_{3}}$ be a tensor. We use $A:=\mathrm{mat}(\mathsf{A})$ to denote a $d_{1}\times(d_{2}\cdot d_{3})$ size matrix. $\mathsf{A}_{a,b,c}=A_{a,(b-1)d_{3}+c}$ for any $a\in[d_{1}],b\in[d_{2}],c\in[d_{3}]$. Let $I_{d}\in\mathbb{R}^{d\times d}$ denote an identity matrix. Let $\mathsf{I}_{d}\in\mathbb{R}^{d\times d\times d}$ denote an identity tensor, i.e., the diagonal entries are $1$ and zeros everywhere else. Let $X\in\mathbb{R}^{d\times d^{2}}$. Let $x\in\mathbb{R}^{d^{3}}$ denote the vectorization of $X\in\mathbb{R}^{d\times d^{2}}$. Let $\mathsf{X}\in\mathbb{R}^{d\times d\times d}$ be the tesnorization of $X\in\mathbb{R}^{d\times d^{2}}$.
@@ 3.1 Definition of Tensor Operations
Here, we define some tensor operations named Kronecker product, which is a matrix operation applied to two matrices of any size, producing a block matrix. It represents a specific type of tensor product and is different from regular matrix multiplication.
Definition 3.1.
We use $\circ$ to denote the Hadamard product i.e. the $(i,j)$-entry of $A\circ B$ is $A_{i,j}B_{i,j}$.
Fact 3.2.
For all $a,b\in\mathbb{R}^{n}$, we have $\mathrm{diag}(a)\cdot b=\mathrm{diag}(b)\cdot a=a\circ b=b\circ a$.
@@ 3.2 Key Definitions of Tensor Attention
Now, we are ready to introduce the tensor attention. First, we introduce the parameters and input for a tensor attention head.
@@
Then, based on the Kronecker product, we can define our tensor attention in the following way.
Definition 3.7 (Tensor attention, Definition 1.1 in [21]).
@@
where
  • $A\in\mathbb{R}^{n\times n^{2}}$ which is defined as $A:=\exp(QK^{\top}/d)$, where $K:=K_{1}\oslash K_{2}\in\mathbb{R}^{n^{2}\times d}$,
  • $D\in\mathbb{R}^{n\times n}$ which is defined as $D:=\operatorname{diag}(A{\bf 1}_{n^{2}})$, and
  • $V\in\mathbb{R}^{n^{2}\times d}$ which is defined as $V:=V_{1}\oslash V_{2}$.
@@
Our Definition 3.7 covers both the cross-attention and self-attention setting. If $Q,K_{1},K_{2},V_{1},V_{2}$ shares the same input, then it is a tensor self-attention, which can capture high-order information. If $Q,K_{1},K_{2},V_{1},V_{2}$ has different modality input, then it is a tensor cross-attention, which can capture high-order relationships among multi-modality.
Also, note that we have $A\in\mathbb{R}^{n\times n^{2}}$ in Definition 3.7. Although $QK^{\top}$ is a low-rank matrix with rank at most $d$, $\exp(QK^{\top})$ may be a full-rank matrix in general. Thus, it is clear to see the exact forward computation of tensor attention takes $\Omega(n^{3})$ time complexity. Here, we introduce a forward tensor attention approximation task, which will help us formulate the tensor attention gradient approximation task later. Furthermore, [21] shows that they can solve this approximation task in almost linear time $n^{1+o(1)}$ (Lemma 5.1).
@@
Given input matrices $Q,K_{1},K_{2},V_{1},V_{2}\in\mathbb{R}^{n\times d}$ and parameters $\epsilon,B>0$, where $\max\{\|Q\|_{\infty},$ $\|K_{1}\|_{\infty},$ $\|K_{2}\|_{\infty},\|V_{1}\|_{\infty},$ $\|V_{2}\|_{\infty}\}\leq B$. Then, our target is to output a matrix $T\in\mathbb{R}^{n\times d}$ satisfying
$\|\underbrace{T}_{n\times d}-\underbrace{D^{-1}}_{n\times n}\underbrace{A}_{n\times n^{2}}\underbrace{V}_{n^{2}\times d}\|_{\infty}\leq\epsilon,$
where $A,D,V$ are defined in Definition 3.7.
Figure 3: The visualization of loss function defined in Definition 3.10. Let $A_{1},A_{2},A_{3},A_{4},A_{5}$ and $E$ be $n\times d$ input matrices. Let $Y$ be a given matrix with size $d^{2}\times d$. The Kronecker product operator $\otimes$ is defined in Definition 3.3. We minimize matrix $X\in\mathbb{R}^{d\times d^{2}}$ in our loss function. The dimensions of all matrices are given in the figure. We first compute $\exp(A_{1}X(A_{2}\otimes A_{3})^{\top})$. Then, we can compute $D(X):=\operatorname{diag}(\exp(A_{1}X(A_{2}\otimes A_{3})^{\top}){\bf 1}_{n^{2}})$. Afterwards, we compute $D(X)^{-1}\exp(A_{1}X(A_{2}\otimes A_{3})^{\top})(A_{4}\otimes A_{5})Y-E$. Finally, we optimize $X$ to compute the minimum of its Frobenius norm with a scaling factor $0.5$.
Similarly, we introduce our tensor attention optimization task. See Figure 3 for more details.
@@
Suppose that $A_{1},A_{2},A_{3},A_{4},A_{5},E\in\mathbb{R}^{n\times d}$ and $Y_{1},Y_{2}\in\mathbb{R}^{d\times d}$ are given. We can formulate the attention optimization problem as
@@
where
  • $A_{2}\otimes A_{3}\in\mathbb{R}^{n^{2}\times d^{2}}$ is the tensor product between $A_{2}$ and $A_{3}$
  • $D(X)=\operatorname{diag}(\exp(A_{1}XA_{2}\otimes A_{3})^{\top}/d){\bf 1}_{n^{2}})\in\mathbb{R}^{n\times n}$
  • $Y=Y_{1}\oslash Y_{2}\in\mathbb{R}^{d^{2}\times d}$
Remark 3.11.
Let weight matrices $W_{Q},W_{K_{1}},W_{K_{2}},W_{V_{1}},W_{V_{2}}\in\mathbb{R}^{d\times d}$ be defined in Definition 3.6. For the self-attention setting, we can see $A_{1},A_{2},A_{3},A_{4},A_{5}\in\mathbb{R}^{n\times d}$ as $Z\in\mathbb{R}^{n\times d}$ in Definition 3.6, and we can see $X\in\mathbb{R}^{d\times d^{2}}$ in Definition 3.10 as $W_{Q}\cdot(W_{K_{1}}\oslash W_{K_{2}})^{\top}\in\mathbb{R}^{d\times d}$ and see $Y\in\mathbb{R}^{d^{2}\times d}$ as $W_{V_{1}}\oslash W_{V_{2}}\in\mathbb{R}^{d^{2}\times d}$. To overcome the cubic complexity obstacle, we only need to handle the gradient computation of $W_{Q}\cdot(W_{K_{1}}\oslash W_{K_{2}})^{\top}\in\mathbb{R}^{d\times d}$, which is in non-linear function $\mathsf{softmax}$.
Although $X\in\mathbb{R}^{d\times d^{2}}$, we have the naive tensor attention gradient computation takes $\Omega(n^{3})$ time, as $A_{1}X(A_{2}\otimes A_{3})^{\top}\in\mathbb{R}^{n\times n^{2}}$. Thus, we formulate the below Approximate Tensor Attention Loss Gradient Computation task as our main focus.

### v2

@@
Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
Tensor Attention, a multi-view attention that is able to capture high-order correlations among multiple modalities, can overcome the representational limitations of classical matrix attention. However, the $O(n^{3})$ time complexity of tensor attention poses a significant obstacle to its utilization in transformers, where $n$ is the input sequence length. In this work, we prove that the backward gradient of tensor attention training can be computed in almost linear time $n^{1+o(1)}$, the same complexity as its forward computation under the bounded entries assumption. We provide a closed-form solution for the gradient and propose a fast computation method utilizing polynomial approximation methods and tensor algebraic techniques. Furthermore, we prove the necessity and tightness of our assumption through hardness analysis, showing that slightly weakening it renders the gradient problem unsolvable in truly subcubic time. Our theoretical results establish the feasibility of efficient higher-order transformer training and may facilitate practical applications of tensor attention architectures.
@@ 1 Introduction
The generative large language models (LLMs), such as Mistral [49], Llama [105], Llama2 [107], Llama3 [10], Gemma [106], GPT-3 [21], GPT-4 [2], Claude3 [12], Grok-1 [118] and many more, have been widely involved in people’s living and work in these two years, such as bio-informatics [108], coding [45], education [54], finance [72], law [100], and even writing NeurIPS conference reviews [59]. The success of LLMs is based on the transformer architecture introduced by [111], which also has been introduced into other modality [30], such as vision-language models, e.g., CLIP [87], Flamingo [7], LLaMA-Adapter [128, 36], LLava [68, 60], BLIP [69, 65], MiniGPT-4 [124], Qwen [17, 18], Gemini [103], MM1 [76].
The above open-sourced large models use two-view matrix attention, i.e., each attention score/entry is related to two tokens (one query token and one key token) to capture the data correlation. More specifically, let $Z$ be hidden representations and $Q=ZW_{Q},K=ZW_{K},V=ZW_{V}$ be the corresponding query, key, and value matrices after projections using weights $W_{Q},W_{K},W_{V}$ respectively. Then, the classical/matrix attention head can be written as $\mathsf{Att}(Z)=\mathsf{Softmax}(QK^{\top})V$.
On the other hand, many studies find that multi-view is crucial for high-order correlation in various kinds of data, e.g., math [96], graph [31, 73], and multi-modality [57]. For example, recently, OpenAI released GPT-4o [83], and Google released Project Astra [37], two flagship multi-modality models that aim to reason across three views, i.e., audio, vision, and text in real-time.
However, [96] theoretically and empirically shows that classical attention can capture pairwise correlation but not triple-wise correlation due to the representational limitations of matrix attention. In other words, one classical matrix attention head “cannot” capture the information relevant to multi-views simultaneously unless using multiple layers with careful architecture design. This poses a fundamental technical obstacle for multi-modality models to efficiently fuse multiple representations/views to capture the high-order correlation among them.
Table 1: Comparison to previous works.
@@
Figure 1: The visualization of tensor attention with $\mathsf{Softmax}$ activation function (Definition 3.5). We give an example of input token length $n=8$, feature dimension $d=4$.
There are several recent works to overcome the time complexity bottleneck above, e.g., $O(n^{2})$ for matrix attention and $O(n^{3})$ for tensor attention. [127] accelerate matrix attention forward via kernel density estimation and get truly sub-quadratic time running time. [13] uses the polynomial approximation method to map the matrix attention into low-rank matrices during forward computation, leading to an almost linear time complexity $n^{1+o(1)}$ when entries are bounded. Similarly, under sparsity assumptions, [41] achieves nearly linear time computation for matrix attention forward by identifying the larger entries in the attention matrix. On the one hand, with fine-grained analysis, [14] proposes a new backward algorithm to compute the gradient of matrix attention in almost linear time complexity $n^{1+o(1)}$ as well, under the same bounded entry assumption. On the other hand, [15] surprisingly finds that the forward computation of tensor attention can also be achieved in almost linear time $n^{1+o(1)}$ rather than almost quadratic time $n^{2+o(1)}$, under similar assumptions as [13]. See a summary in Table 1. Thus, it is natural to ask,
Can we achieve almost linear time for gradient computation in Tensor Attention Training?
We provide a positive answer in this work. Under the same bounded entries assumption as [15], we propose Algorithm 1 to fast compute the backward gradient of Tensor Attention Training in almost linear time $n^{1+o(1)}$ as its forward computation. Thus, our results may make the tensor attention practical, as we can get around the $O(n^{3})$ complexity barrier both in its forward and backward computation. Our contributions are summarized as follows:
@@
  • Based on the closed-form solution, by utilizing polynomial approximation methods and tensor computation techniques, we propose Algorithm 1 to fast compute the backward gradient of tensor attention training in almost linear time as its forward computation (Theorem 5.2).
@@ 2 Related Work
In recent years, significant advances have been made in the development of efficient attention computation. One research direction involves employing low-rank approximations, polynomial kernel, or random features for the attention matrix [130, 1, 13, 52, 95, 4, 44, 113, 39, 40, 70, 102, 38], which scales the computational complexity sub-quadratically with sequence length. Another method explores patterns of sparse attention that lessen the computational load [93, 41, 63, 67, 64, 62]. Additionally, using linear attention as an alternative to softmax attention has emerged as a substantial area of study [104, 55, 97, 126, 94, 4, 101, 123, 66, 33]. These innovations have enhanced the capability of transformer-based models to handle longer sequences, thereby broadening their potential applications across various fields [27, 91, 86, 34, 80, 120, 9, 115, 16, 71, 23, 24, 47, 98].
@@
Tensors excel over matrices in capturing higher-order relationships within data. Calculating low-rank factorizations or approximations of tensors is essential in a wide range of computer science applications, such as natural language processing [28, 74, 75, 22], computer vision [8, 85, 58, 25], computer graphics [112, 117, 110], security [3, 6, 51], and data mining [50, 89, 53, 79]. Moreover, tensors are crucial in numerous machine learning applications [84, 48, 42, 11, 129, 122, 99] and other diverse fields [90, 26, 125, 121, 88].
Roadmap. In Section 3, we introduce the notations, several useful definitions, and our loss function. In Section 4, we give the closed form of the gradient of our loss function, and also its computational time complexity. In Section 5, we prove that we can compute the gradient in almost linear time. In Section 6, we provide the hardness analysis. In Section 7, we give the conclusion of our paper.
@@ 3 Preliminary
In this section, we first provide the notations we use. In Section 3.1, we provide general definitions related to tensor operation. In Section 3.2, we provide key definitions that we utilize in this paper.
Basic notations. We use $[n]$ to denote $\{1,2,\dots,n\}$. We use $e_{i}$ to denote a column vector where only $i$-th location is $1$ and zeros everywhere else. We denote an all $1$ vector using ${\bf 1}_{n}\in\mathbb{R}^{n}$ We use $\langle a,b\rangle$ to denote the inner product of $a,b\in\mathbb{R}^{d}$ i.e. $\langle a,b\rangle:=\sum_{i=1}^{d}a_{i}b_{i}$. We use $\|x\|_{p}$ to denote the $\ell_{p}$ norm of a vector $x\in\mathbb{R}^{n}$, i.e. $\|x\|_{2}:=(\sum_{i=1}^{n}x_{i}^{2})^{1/2}$, and $\|x\|_{\infty}:=\max_{i\in[n]}|x_{i}|$. We use $\circ$ to denote the Hadamard product i.e. the $(i,j)$-entry of $A\circ B$ is $A_{i,j}B_{i,j}$.
We use $\operatorname{tr}[A]:=\sum_{i=1}^{n}A_{i,i}$ to denote the trace of a matrix $A\in\mathbb{R}^{n\times n}$. We use $\exp(A)$ to denote a matrix where $\exp(A)_{i,j}:=\exp(A_{i,j})$ for a matrix $A\in\mathbb{R}^{n\times d}$. We use $\|A\|_{\infty}$ to denote the $\ell_{\infty}$ norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e. $\|A\|_{\infty}:=\max_{i\in[n],j\in[d]}|A_{i,j}|$. We use $\|A\|_{F}$ to denote the Frobenius norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e. $\|A\|_{F}:=\sqrt{\sum_{i\in[n]}\sum_{j\in[d]}|A_{i,j}|^{2}}$. We use $\operatorname{poly}(n)$ to denote polynomial time complexity w.r.t $n$.
Tensor related notations.
Let $A\in\mathbb{R}^{n\times d}$. We use $a:=\operatorname{vec}(A)$ to denote a length $nd$ vector. We stack rows of $A$ into a column vector, i.e. $\operatorname{vec}(A):=[a_{1}^{\top},a_{2}^{\top},\dots,a_{n}^{\top}]^{\top}$ where $a_{i}^{\top}$ is the $i$-th row of $A$, or simply $\operatorname{vec}(A)_{j+(i-1)d}:=A_{i,j}$ for any $i\in[n],j\in[d]$, visualized in Fig. 2. Let $I_{d}\in\mathbb{R}^{d\times d}$ denote an identity matrix. Let $\mathsf{I}_{d}\in\mathbb{R}^{d\times d\times d}$ denote an identity tensor, i.e., the diagonal entries are $1$ and zeros everywhere else. Let $X\in\mathbb{R}^{d\times d^{2}}$. Let $x\in\mathbb{R}^{d^{3}}$ denote the vectorization of $X\in\mathbb{R}^{d\times d^{2}}$. Let $\mathsf{X}\in\mathbb{R}^{d\times d\times d}$ be the tesnorization of $X\in\mathbb{R}^{d\times d^{2}}$, where $\mathsf{X}_{a,b,c}=X_{a,(b-1)d+c}$ for any $a,b,c\in[d]$. And we have $X=\mathrm{mat}(\mathsf{X})$.
@@ 3.1 Definition of Tensor Operations
We define some operations like the Kronecker product, which is a matrix operation applied to two matrices of any size, producing a block matrix. It is different from regular matrix multiplication and will be useful for our introduction and analysis of tensor attention.
@@ 3.2 Key Definitions of Tensor Attention
Now, we are ready to introduce the tensor attention. First, we introduce the parameters and input.
@@
Then, based on Kronecker product, we define tensor attention in the following way.
Definition 3.5 (Tensor attention, Definition 7 in [96], Definition 1.1 in [15]).
@@
where (1) $A:=\exp(QK^{\top}/d)\in\mathbb{R}^{n\times n^{2}}$ and $K:=K_{1}\oslash K_{2}\in\mathbb{R}^{n^{2}\times d}$, (2) $D:=\operatorname{diag}(A{\bf 1}_{n^{2}})\in\mathbb{R}^{n\times n}$, and (3) $V:=V_{1}\oslash V_{2}\in\mathbb{R}^{n^{2}\times d}$.
@@
Our Definition 3.5 covers the self-attention setting, when the query/key/values $Q,K_{1},K_{2},V_{1},V_{2}$ follow Definition 3.4 where they share the same input. It is then a tensor self-attention, which can capture high-order information of the input $Z$. When the query/key/values have different inputs, it is then a tensor cross-attention that can capture high-order relationships among multiple inputs.
Also, note that we have $A\in\mathbb{R}^{n\times n^{2}}$ in Definition 3.5. Although $QK^{\top}$ is a low-rank matrix with rank at most $d$, $\exp(QK^{\top})$ may be a full-rank matrix in general. Thus, it is clear to see the exact forward computation of tensor attention takes $O(n^{3})$ time. Here, we introduce a forward tensor attention approximation task, which will help us formulate the tensor attention gradient approximation task later. Furthermore, [15] show that they can solve this approximation task in almost linear time $n^{1+o(1)}$ (Lemma 5.1).
@@
Given input matrices $Q,K_{1},K_{2},V_{1},V_{2}\in\mathbb{R}^{n\times d}$ and parameters $\epsilon,B>0$, where $\max\{\|Q\|_{\infty},$ $\|K_{1}\|_{\infty},$ $\|K_{2}\|_{\infty},\|V_{1}\|_{\infty},$ $\|V_{2}\|_{\infty}\}\leq B$. Let $A,D,V$ be defined in Definition 3.5. Then, our target is to output a matrix $T\in\mathbb{R}^{n\times d}$ satisfying
$\|\underbrace{T}_{n\times d}-\underbrace{D^{-1}}_{n\times n}\underbrace{A}_{n\times n^{2}}\underbrace{V}_{n^{2}\times d}\|_{\infty}\leq\epsilon.$
Figure 3: The visualization of loss function defined in Definition 3.8. Let $A_{1},A_{2},A_{3},A_{4},A_{5}$ and $E$ be $n\times d$ input matrices. Let $Y$ be a given matrix with size $d^{2}\times d$. The Kronecker product operator $\otimes$ is defined in Definition 3.1. We minimize matrix $X\in\mathbb{R}^{d\times d^{2}}$ in our loss function. We first compute $\exp(A_{1}X(A_{2}\otimes A_{3})^{\top})$. Then, we compute $D(X):=\operatorname{diag}(\exp(A_{1}X(A_{2}\otimes A_{3})^{\top}){\bf 1}_{n^{2}})$. Afterwards, we compute $D(X)^{-1}\exp(A_{1}X(A_{2}\otimes A_{3})^{\top})(A_{4}\otimes A_{5})Y-E$. Finally, we optimize $X$ to compute the minimum of its Frobenius norm with a scaling factor $0.5$.
For our focus, tensor attention training, we would like to find weights to fit the tensor attention to a desired output $E$. We first simplify the attention expression of Definition 3.5, whose inputs are from Definition 3.4 with weight matrices $W_{Q},W_{K_{1}},W_{K_{2}},W_{V_{1}},W_{V_{2}}\in\mathbb{R}^{d\times d}$. Let $X:=W_{Q}\cdot(W_{K_{1}}\oslash W_{K_{2}})^{\top}\in\mathbb{R}^{d\times d^{2}}$ and $Y:=W_{V_{1}}\oslash W_{V_{2}}\in\mathbb{R}^{d^{2}\times d}$. It can be verified that the tensor attention equals $D^{-1}\exp(ZX(Z\otimes Z)^{\top}/d)(Z\otimes Z)Y$. Then we model the tensor attention training as the following tensor attention optimization problem (where $A_{1},A_{2},A_{3},A_{4},A_{5}$ are introduced to replace $Z$ to capture more general settings such as cross-attention). See Figure 3 for an illustration.
@@
Suppose that $A_{1},A_{2},A_{3},A_{4},A_{5},E\in\mathbb{R}^{n\times d}$ and $Y_{1},Y_{2}\in\mathbb{R}^{d\times d}$ are given. We formulate the attention optimization problem as
@@
where (1) $A_{2}\otimes A_{3}\in\mathbb{R}^{n^{2}\times d^{2}}$ is the tensor product between $A_{2}$ and $A_{3}$, (2) $D(X)=\operatorname{diag}(\exp(A_{1}X(A_{2}\otimes A_{3})^{\top}/d){\bf 1}_{n^{2}})\in\mathbb{R}^{n\times n}$, and (3) $Y=Y_{1}\oslash Y_{2}\in\mathbb{R}^{d^{2}\times d}$.
The naive gradient computation for the above tensor attention training takes $\Omega(n^{3})$ time. The gradient for $X$ is the bottleneck while that for $Y$ is not, since $A_{1}X(A_{2}\otimes A_{3})^{\top}\in\mathbb{R}^{n\times n^{2}}$ lies in the non-linear function $\mathsf{Softmax}$. Also, note that with gradients of $X$ and $Y$, it is easy to get the gradients of the weight matrices $W_{Q},W_{K_{1}},W_{K_{2}},W_{V_{1}},W_{V_{2}}$. Therefore, our main focus is the following Approximate Tensor Attention Loss Gradient Computation task.